# Quantitative Propagation of Chaos for 2D Viscous Vortex Model with General Circulations on the Whole Space

### v1

@@ 1.2. Main Results
The main result of this paper is to provide the quantitative propagation of chaos results of the point vortex system (1.1) on the 2D whole space. We shall both work in the global and local sense. In this paper, the global convergence is termed as the asymptotic vanishing of $H_{N}(G^{N}|g^{N})$, while the local convergence is for the asymptotic vanishing of $H_{k}(G^{N,k}|g^{k})$.
For the global case, inspired by the groundbreaking paper [JW18], we consider the global normalized relative entropy between the joint law of the particle system and the tensorized law of the limit system, which controls the $L^{1}$ distance of marginals between two systems by the Csiszár-Kullback-Pinsker (CKP) inequality. By estimating the time-derivative of the relative entropy, using the Law of Large Numbers and Large Deviation type results introduced by [JW18], we are able to close the Grönwall loop and arrive at a global optimal decay rate, hence proving the strong propagation of chaos result. We mention that in this case we prove our results for all positive viscosity $\sigma>0$. This part of our main results is an extension of the previous work [FHM14] for qualitative propagation of chaos and the previous work by the authors [FW23] for the special case ${\mathcal{M}}_{j}\equiv 1$ and by Shao-Zhao [SZ24] for the torus case.
@@
Assume that $G^{N}\in\mathcal{P}({\mathbb{R}}^{N}\times{\mathbb{R}}^{2N})$ is an entropy solution of the Liouville equation (1.5) and that $g\in L^{\infty}([0,T];W^{2,1}\cap W^{2,\infty}({\mathbb{R}}\times{\mathbb{R}}^{2}))$ solves (1.7) with $g\geq 0$ and ×$\int_{{\mathbb{R}}\times{\mathbb{R}}^{2}}g(z)\,\mathrm{d}z=1$. Assume further that the initial data $g_{0}\in W_{x}^{2,1}\cap W_{x}^{2,\infty}({\mathbb{R}}\times{\mathbb{R}}^{2})$ satisfies the logarithmic growth conditions
@@
Applying the Csiszár-Kullback-Pinsker inequality and using that $A\leq 1$ , we establish the global $L^{1}$ strong propagation of chaos towards the solution to the vorticity formulation of the 2D Navier-Stokes equation (1.4).
@@
Under the same assumption as Theorem 1.1, define the $k$-particle vorticity $\omega^{N,k}$ by
$\omega^{N,k}(x_{1},...,x_{k})=\int_{{\mathbb{R}}^{k}\times({\mathbb{R}}\times{\mathbb{R}}^{2})^{N-k}}m_{1}...m_{k}G^{N}(Z^{N})\,\mathrm{d}m_{1}...\,\mathrm{d}m_{k}\,\mathrm{d}z_{k+1}...\,\mathrm{d}z_{N}.$
@@
$\|\omega^{N,k}-\omega^{\otimes k}\|_{TV}\leq\|G^{N,k}-g^{\otimes k}\|_{TV}\leq\frac{\sqrt{k}}{\sqrt{N}}Me^{M\log^{2}(1+t)}$
@@
For the local case, as observed firstly by Lacker [Lac23], in order to obtain the optimal convergence rate in $N$ between $G^{N,k}$ and $g^{k}$ for any fixed $k$, we shall consider the local normalized relative entropy directly instead of using the sub-additivity property
@@
as shown in Corollary 1.1 above. However, the best optimal convergence rate one may expect is actually $O(k/N)$, tested by some simple Gaussian examples. Using the BBGKY hierarchy to construct an ODE hierarchy of relative entropy $H_{k}$ of all orders, this optimal rate has been verified for Lipschitz kernels in [Lac23]. The recent work by S. Wang [Wan24] has extended this idea to the 2D Navier-Stokes equation with fixed circulations ${\mathcal{M}}_{i}\equiv 1$ on the torus, being the first attempt for solving the singular kernel case. Our results shall improve it to the general circulations and to the whole space, but we mention that our result is still restricted to the high viscosity case, say large enough $\sigma>0$, as in [Wan24].
@@
$\sigma>\sqrt{2}A\|V\|_{\infty}.$
Then we have
$H_{k}(G^{N,k}|g^{k})\leq Me^{M\log^{2}(1+t)}\Big{(}H_{k}(G_{0}^{N,k}|g_{0}^{k})+\frac{k}{N^{2}}\Big{)}$ (1.13)
@@
Applying again the Csiszár-Kullback-Pinsker inequality and using that $A\leq 1$ , we establish the sharp local $L^{1}$ strong propagation of chaos towards the solution to the vorticity formulation of the 2D Navier-Stokes equation (1.4).
@@
$\|\omega^{N,k}-\omega^{\otimes k}\|_{TV}\leq\frac{k}{N}Me^{M\log^{2}(1+t)}$
@@
We should state here why we cannot assume that ${\mathcal{M}}_{i}$ and $X_{i}(0)$ are mutually independent in our vortex model. This is essentially required when studying propagation of chaos results for this system, since we need the Kac’s chaos condition on the initial data, say
@@
${\mathbb{E}}\Big{\langle}\frac{1}{N}\sum_{i=1}^{N}{\mathcal{M}}_{i}\delta_{X_{i}(0)},\varphi\Big{\rangle}={\mathbb{E}}{\mathcal{M}}\cdot{\mathbb{E}}\varphi\Big{(}X(0)\Big{)}\rightarrow\int_{{\mathbb{R}}^{2}}\varphi(x)\omega_{0}(x)\,\mathrm{d}x.$
@@
where $f_{0}(x)$ is the probability density of $X(0)$. This yields that $\omega_{0}(x)$, and hence $\omega_{t}(x)$ by parabolic maximum principle, must be always non-positive or non-negative, depending on whether ${\mathbb{E}}{\mathcal{M}}$ is non-positive or not. Hence we will not able to approximate general vorticity which may change sign. This will eliminate the physical meaning of our study, since we expect that our results can provide particle approximation arguments to general vorticity by constructing the system with general circulations. This coincides with our previous remark about the significance of the appearance of the circulation ${\mathcal{M}}_{j}$ in the point vortex system (1.1). One may encounter with the same restriction if the initial density of the limit system is assumed to be separating variables, say
@@
although this will provide convenience in mathematical calculations. Hence we also cannot make this assumption as done in [SZ24].
Nevertheless, we emphasize that the independence assumption is not necessary for dealing with the torus case in [SZ24], though the authors do not point it out in their paper. For completeness, we will give a short sketch of the proof of the torus case without independence assumption in our proof of main results, but this result is essentially due to [SZ24] and we shall claim no originality to this.
@@ 1.3. Related Works
Propagation of chaos and mean-field limit for large systems of interacting particles, widely appearing in analysis of PDEs and probability theory, has been extensively studied over the last decades, both for the first-order and second-order systems. The derivation of some effective equations to describe the average or statistical behavior of large interacting particle systems can be traced back to Maxwell and Boltzmann. Kac [Kac56] first raises the basic notions of chaos and proves some partial propagation of chaos results for the simplified model of 1D Boltzmann equation, and now the related studies have been far beyond the range of kinetic theory. The fundamental mathematical setting of this problem is set up and basically resolved by McKean [McK67] and Dobrushin [Dob79] using the classical coupling method. Since then people began to put attention to the much more difficult models, say systems with singular interacting kernels, instead of the regular (bounded Lipschitz) kernels. As for our point vortex model approximating the famous 2D Navier-Stokes/Euler equation, the first result dates back to Osada [Osa86] with bounded initial data and large viscosity. Later Fournier-Hauray-Mischler [FHM14] extend the result to all positive viscosity cases, while the initial data is only assumed to have finite moments and entropy, using the compactness argument and the entropy estimate.
@@
The relative entropy method as in Jabin-Wang [JW16, JW18] has seen great effectiveness in solving quantitative propagation of chaos results for McKean-Vlasov process in recent years, especially making sense for the singular interacting kernel cases. In [JW16], Jabin and the 2nd author resolved the second-order case with bounded kernels. In [JW18] they successfully applied to the 2D Navier-Stokes equation with the singular kernel given by the Biot-Savart law, but restricted to the torus case and without general circulations. Later extended results include Wynter [Wyn21] for the mixed-sign case, Guillin-Le Bris-Monmarché [GLBM24] for uniform-in-time propagation of chaos using the logarithmic Sobolev inequality (LSI) for the limit density, and Shao-Zhao [SZ24] for general circulations with common environmental noise, all of which are still working on the torus. In our previous work [FW23] we extended the result to the whole Euclidean space, given some special growth conditions on the initial data, which are natural and reasonable since those estimates automatically hold for general $L^{1}$ and Gaussian decaying initial data as long as the time becomes positive. These arguments appear to be more reasonable when considering the strengthening chaos results for instance by Huang [Hua23], where the entropic propagation of chaos is derived for any positive time given only initial chaos in Wasserstein sense, for bounded interacting kernels. We also mention the recent applications of the relative entropy method to establish propagation of chaos for other models, for instance the Landau-Maxwellian equation case by Carrillo-Feng-Guo-Jabin-Wang [CFG+24] and some uniform-in-time results for second-order systems with regular enough kernels by Gong-Wang-Xie [GWX24].
@@
The modulated energy method introduced by Serfaty [Ser20] is another effective tool for resolving propagation of chaos results for singular kernel case, especially the Coulomb/Riesz case. By constructing a Coulomb/Riesz based metric between the empirical measure of the $N$-particle system and the limit density. Serfaty [Ser20] established the quantitative propagation of chaos for the first-order Coulomb and super-Coulomb cases without noises in the whole space, in the sense of controlling the weak convergence rate when acting on test functions. This method has been extended to various problems, including Rosenzweig [Ros22] for 2D incompressible Euler equation, Nguyen-Rosenzweig-Serfaty [NRS22] for Riesz-type flows, Rosenzweig-Serfaty [RS23a] for uniform-in-time estimates for sub-Coulomb cases, De Courcel-Rosenzweig-Serfaty [dCRS23b] for periodic Riesz cases. The key estimates appearing in the control of modulated energy, say the commutator type estimates, have been thoroughly examined in the recent work Rosezweig-Serfaty [RS24b]. The combination of modulated energy and relative entropy leads to the modulated free energy method, which is applied to the Coulomb/Riesz kernels on the torus and more interestingly the first time to the attractive kernel cases by Bresch-Jabin-Wang [BJW19b, BJW19a, BJW23] for the 2D Patlak-Keller-Segel model. See also the recent work by De Courcel-Rosenzweig-Serfaty [dCRS23a] for further results on uniform-in-time propagation of chaos and phase transition for attractive log gas. In the forthcoming work Cai-Feng-Gong-Wang [CFGW24], the authors and coauthors are also able to apply the modulated free energy method to resolve the 2D Coulomb case on the whole space. Some generation of chaos results are also obtained partially by Rosenzweig-Serfaty [RS23b, RS24a] using the time-uniform LSI for the limit density or the modulated Gibbs measure. We refer to the recent lecture notes by Serfaty [Ser24] for more complete discussions on the Coulomb/Riesz gases concerning above topics.
@@
We also emphasize that in the work of Lacker [Lac23] and Lacker-Le Flem [LLF23], it is shown that using the BBGKY hierarchy that the optimal local convergence rate of the relative entropy between $k$-marginals is $O(k/N)$ for systems with bounded interactions, while previous global arguments can only obtain the suboptimal $O(\sqrt{k/N})$ rate. This is verified by calculating the local relative entropy and solving the ODE hierarchy by iteration. Most recently in S. Wang [Wan24] the optimal rate has been extended to the 2D Biot-Savart case with large viscosity on the torus also using the BBGKY hierarchy. Qualitative propagation of chaos for the 2D Vlasov-Poisson equation has also been established by Bresch-Jabin-Soler [BJS22] by hierarchy estimates. Recently, a new duality based method for the mean field limit problem was introduced by Bresch-Duerinckx-Jabin [BDJ24], which requires only that the interacting kernel is in $L^{2}$.
@@ 1.4. Outline of the Article
The rest of this paper is organized as follows. We first present the proof of our main results in Section 2, acknowledging some logarithmic growth estimates of the limit density. The proof is divided into several subsections, respectively giving the time evolution of the relative entropy, some useful functional inequality lemmas and theorems, and the main controlling arguments. We shall provide the global and local relative entropy control respectively, and also set up an independent subsection for the proof of an ODE hierarchy proposition. In Section 3 we apply the Carlen-Loss arguments and the parabolic maximum principle to give the needed regularity estimates for the limit density. In Section 4 we first present the Grigor’yan parabolic maximum principle as the key tool, and then provide our desired logarithmic growth estimates, i.e. growth controls on $\nabla\log g$ and $\nabla^{2}\log g$ under certain initial conditions.
@@ 2.1. Time Evolution of Relative Entropy
We present in this subsection the time evolution of the normalized relative entropy, which is computed directly using the evolution PDE of the two systems. First we give the calculations for the global relative entropy.
@@
$\displaystyle\frac{\,\mathrm{d}}{\,\mathrm{d}t}H_{N}(G^{N}|g^{N})=$ $\displaystyle-\frac{\sigma}{N}\sum_{i=1}^{N}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{N}}G^{N}\Big{|}\nabla_{x_{i}}\log\frac{G^{N}}{g^{N}}\Big{|}^{2}\,\mathrm{d}Z^{N}$
$\displaystyle-\frac{1}{N^{2}}\sum_{i,j=1}^{N}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{N}}G^{N}\Big{(}m_{j}K(x_{i}-x_{j})-K\ast\omega(x_{i})\Big{)}\cdot\nabla\log g(m_{i},x_{i})\,\mathrm{d}Z^{N}.$
@@
while the tensorized law $g^{N}$ solves (1.8):
@@
$\displaystyle$=\,= $\displaystyle\frac{1}{N}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{N}}\partial_{t}G^{N}(\log G^{N}+1-\log g^{N})+\frac{G^{N}}{g^{N}}\partial_{t}g^{N}\,\mathrm{d}Z^{N}$
$\displaystyle$=\,= $\displaystyle\frac{1}{N}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{N}}\Big{(}\sigma\sum_{i=1}^{N}\Delta_{x_{i}}G^{N}-\frac{1}{N}\sum_{i,j=1}^{N}m_{j}K(x_{i}-x_{j})\cdot\nabla_{x_{i}}G^{N}\Big{)}(\log G^{N}+1-\log g^{N})\,\mathrm{d}Z^{N}$
$\displaystyle+\frac{1}{N}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{N}}\Big{(}\sigma\sum_{i=1}^{N}\Delta_{x_{i}}g^{N}-\sum_{i=1}^{N}K\ast\omega(x_{i})\cdot\nabla_{x_{i}}g^{N}\Big{)}\frac{G^{N}}{g^{N}}\,\mathrm{d}Z^{N}.$
@@
$\displaystyle-\frac{\sigma}{N}\sum_{i=1}^{N}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{N}}\nabla_{x_{i}}G^{N}\cdot\nabla_{x_{i}}(\log G^{N}-\log g^{N})\,\mathrm{d}Z^{N}$
$\displaystyle+\frac{1}{N^{2}}\sum_{i,j=1}^{N}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{N}}m_{j}K(x_{i}-x_{j})\cdot\nabla_{x_{i}}(\log G^{N}-\log g^{N})G^{N}\,\mathrm{d}Z^{N}$
$\displaystyle-\frac{\sigma}{N}\sum_{i=1}^{N}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{N}}\nabla_{x_{i}}g^{N}\cdot\nabla_{x_{i}}\frac{G^{N}}{g^{N}}\,\mathrm{d}Z^{N}$
$\displaystyle-\frac{1}{N}\sum_{i=1}^{N}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{N}}\Big{(}K\ast\omega(x_{i})\cdot\nabla_{x_{i}}g^{N}\Big{)}G^{N}\,\mathrm{d}Z^{N}.$
@@
Now we turn to the calculations for the local relative entropy.

### v2

@@ 1.2. Main Results
The main result of this article is to prove quantitative propagation of chaos for the point vortex system (1.1) on the whole space. We shall both work in the global and local sense. In this article, global convergence is termed as the asymptotic vanishing of $H_{N}(G^{N}|g^{N})$, while local convergence is for the asymptotic vanishing of $H_{k}(G^{N,k}|g^{k})$.
For the global case, inspired by the groundbreaking paper [30], we consider the global normalized relative entropy between the joint law of the particle system and the factorized law of the limit system, which controls the $L^{1}$ distance of marginals between two systems by the Csiszár-Kullback-Pinsker (CKP) inequality. By estimating the time evolution of the relative entropy, using the Law of Large Numbers and Large Deviation type results introduced by [30], we are able to close the Grönwall loop and arrive at a global optimal decay rate, hence proving the strong propagation of chaos result. We mention that in this case we prove our results for all positive viscosity $\sigma>0$. This part of our main results is an extension of the previous work [17] for qualitative propagation of chaos and the previous work by the authors [16] for the special case ${\mathcal{M}}_{j}\equiv 1$ and by Shao-Zhao [52] for the torus case.
@@
Assume that $G^{N}\in\mathcal{P}(\mathbb{D}^{N})$ is an entropy solution of the Liouville equation (1.5) and that $g\in L^{\infty}([0,T];W^{2,1}\cap W^{2,\infty}(\mathbb{D}))$ solves (1.7) with $g\geq 0$ and $\int_{\mathbb{D}}g(z)\,\mathrm{d}z=1$. Assume further that the initial data $g_{0}\in W_{x}^{2,1}\cap W_{x}^{2,\infty}(\mathbb{D})$ satisfies the logarithmic growth conditions
@@
Applying the CKP inequality and using that $A\leq 1$ , we establish the global $L^{1}$ strong propagation of chaos towards the solution to the vorticity formulation of the 2D Navier-Stokes equation (1.4).
@@
Under the same assumption as Theorem 1.1, define the $k$-particle vorticity function $\omega^{N,k}$ by
$\omega^{N,k}(x_{1},\cdots,x_{k})=\int_{{\mathbb{R}}^{k}\times\mathbb{D}^{N-k}}m_{1}\cdots m_{k}G^{N}(Z^{N})\,\mathrm{d}m_{1}\cdots\,\mathrm{d}m_{k}\,\mathrm{d}z_{k+1}\cdots\,\mathrm{d}z_{N}.$
@@
$\|\omega^{N,k}-\omega^{\otimes k}\|_{TV}\leq\|G^{N,k}-g^{\otimes k}\|_{TV}\leq\frac{\sqrt{k}}{\sqrt{N}}M(1+t)^{M}$
@@
For the local case, as observed first by Lacker [34], in order to obtain the optimal convergence rate in $N$ between $G^{N,k}$ and $g^{k}$ for any fixed $k$, we shall consider the local normalized relative entropy directly instead of using the sub-additivity property
@@
as shown in Corollary 1.1 above. However, the best optimal convergence rate one may expect is actually $O(k/N)$, tested by some simple Gaussian examples. Using the BBGKY hierarchy to construct an ODE hierarchy of relative entropy $H_{k}$ of all orders, this optimal rate has been verified for Lipschitz kernels in [34, 35]. The recent work by S. Wang [55] has extended this idea to the 2D Navier-Stokes equation with fixed circulations ${\mathcal{M}}_{i}\equiv 1$ on the torus, being the first attempt for solving the singular kernel case. Our results shall improve it to the general circulations and to the whole space, but we mention that our result is still restricted to the high viscosity case, say large enough $\sigma>0$, as in [55].
@@
$\sigma>\sqrt{2}A\|V\|_{L^{\infty}},$
with $V(x)=\Big{(}-\frac{1}{2\pi}\arctan\frac{x_{1}}{x_{2}}\Big{)}\mathrm{Id}$ satisfying $K=\nabla\cdot V$. Then we have
$H_{k}(G^{N,k}|g^{k})\leq M(1+t)^{M}\Big{(}H_{k}(G_{0}^{N,k}|g_{0}^{k})+\frac{k}{N^{2}}\Big{)}$ (1.13)
@@
Applying again the CKP inequality and using that $A\leq 1$ , we establish the sharp local $L^{1}$ strong propagation of chaos towards the solution to the vorticity formulation of the 2D Navier-Stokes equation (1.4).
@@
$\|\omega^{N,k}-\omega^{\otimes k}\|_{TV}\leq\frac{k}{N}M(1+t)^{M}$
@@
We should state here why we cannot assume that ${\mathcal{M}}_{i}$ and $X_{i}(0)$ are mutually independent in our vortex model (1.1). This is essentially required when studying propagation of chaos for this system, since we need to impose Kac’s chaos condition on the initial data, say
@@
${\mathbb{E}}\Big{\langle}\frac{1}{N}\sum_{i=1}^{N}{\mathcal{M}}_{i}\delta_{X_{i}(0)},\varphi\Big{\rangle}={\mathbb{E}}{\mathcal{M}}\cdot{\mathbb{E}}\varphi(X_{0})\rightarrow\int_{{\mathbb{R}}^{2}}\varphi(x)\,\omega_{0}(x)\,\mathrm{d}x.$
@@
where $f_{0}(x)$ is the probability density of $X(0)$. This shows that $\omega_{0}(x)$, and hence $\omega_{t}(x)$ by the parabolic maximum principle, must be always non-positive or non-negative, depending on whether ${\mathbb{E}}{\mathcal{M}}$ is non-positive or not. Hence we will not be able to approximate general vorticity which may change sign. This will eliminate the physical meaning of our study, as we expect that our results can provide particle approximation arguments to general vorticity functions $\omega_{t}$ by constructing the system with general circulations. This coincides with our previous remark about the significance of the appearance of the circulation ${\mathcal{M}}_{j}$ in the point vortex system (1.1). One may encounter the same restriction if the initial density of the limit system is assumed to be separating variables, say
@@
although this will provide convenience in mathematical computations. Therefore, we also cannot make this assumption as in [52].
However, we also emphasize that the independence assumption is not necessary to deal with the torus case in [52], although they did not point it out in their paper. For completeness, we will give a short sketch of the proof of the torus case without the independence assumption in our proof of main results, but this result is essentially due to [52] and we shall claim no originality to this.
@@ 1.3. Related Works
Propagation of chaos and convergence to mean-field limit for large systems of interacting particles, widely appearing in the analysis of PDEs and probability theory, have been extensively studied over the last decades, both for the first-order and second-order systems. The derivation of some effective equations to describe the average or statistical behavior of large interacting particle systems can be traced back to Maxwell [40] and Boltzmann [2]. Kac [31] first raised the basic notions of chaos and proved some partial propagation of chaos results for the simplified model of the Boltzmann equation in dimension 11 11. Related studies have been far beyond the range of kinetic theory. The fundamental mathematical setting of this problem was set up and basically resolved by McKean [41] and Dobrushin [14] using the classical coupling method. Since then people began to pay attention to the much more difficult models, say systems with singular interacting kernels, instead of the regular (bounded Lipschitz) kernels. As for our point vortex model approximating the famous 2D Navier-Stokes/Euler equation, the first result dates back to Osada [43] with bounded initial data and large viscosity. Later Fournier-Hauray-Mischler [17] extended the result to all positive viscosity cases, while the initial data is only assumed to have finite moments and entropy, using the compactness argument and the entropy estimate.
@@
The relative entropy method as in Jabin-Wang [28, 30] has seen great effectiveness in proving quantitative propagation of chaos results for McKean-Vlasov processes in recent years, especially making sense for the singular interacting kernel cases. In [28], Jabin and the second author resolved the second-order case with bounded kernels. In [30] they successfully applied to the 2D Navier-Stokes equation with the singular kernel given by the Biot-Savart law, but restricted to the torus case and without general circulations. Later extended results include Wynter [56] for the mixed-sign case, Guillin-Le Bris-Monmarché [22] for uniform-in-time propagation of chaos using the logarithmic Sobolev inequality (LSI) for the limit density, and Shao-Zhao [52] for general circulations with common environmental noise, all of which are still working on the torus. In our previous work [16] we extended the result to the whole Euclidean space, given some special growth conditions on the initial data, which are natural and reasonable since those estimates automatically hold for general $L^{1}$ initial data with Gaussian upper bound as long as the time becomes positive. These arguments appear to be more reasonable when considering the strengthening chaos results for instance by Huang [26], where the entropic propagation of chaos is derived for any positive time given only initial chaos in the Wasserstein sense, for bounded interacting kernels. We also mention the recent applications of the relative entropy method to establish propagation of chaos for other models, for instance the Landau equation with Maxwellian molecules by Carrillo-Feng-Guo-Jabin-Wang [11] and some uniform-in-time results for second-order systems with regular enough kernels by Gong-Wang-Xie [19].
@@
The modulated energy method introduced by Serfaty [50] is another effective tool for resolving propagation of chaos for the singular kernel case, especially the Coulomb/Riesz cases. By constructing a Coulomb/Riesz based metric between the empirical measure of the $N$-particle system and the limit density. Serfaty [50] established the quantitative propagation of chaos for the first-order Coulomb and super-Coulomb cases without noise in the whole space, in the sense of controlling the weak convergence rate when acting on test functions. This method has been extended to various problems, including Rosenzweig [45] for 2D incompressible Euler equation, Nguyen-Rosenzweig-Serfaty [42] for Riesz-type flows, Rosenzweig-Serfaty [46] for uniform-in-time estimates for sub-Coulomb cases, and De Courcel-Rosenzweig-Serfaty [13] for periodic Riesz cases. The key estimates appearing in the control of modulated energy, say the commutator type estimates, have been thoroughly examined in the recent work Rosezweig-Serfaty [49]. The combination of modulated energy and relative entropy leads to the modulated free energy method, which is applied to the Coulomb/Riesz kernels on the torus and more interestingly the first time to the attractive kernel cases by Bresch-Jabin-Wang [6, 5, 7] for the 2D log gas and the 2D Patlak-Keller-Segel model. See also the recent work by De Courcel-Rosenzweig-Serfaty [12] for further results on uniform-in-time propagation of chaos and phase transition for attractive log gas. In the recent work Cai-Feng-Gong-Wang [9], the authors and coauthors are also able to apply the modulated free energy method to resolve the 2D Coulomb case on the whole space. Some generation of chaos results are also obtained partially by Rosenzweig-Serfaty [47, 48] using the time-uniform LSI for the limit density or the modulated Gibbs measure. We refer to the recent lecture notes by Serfaty [51] for more complete discussions on the Coulomb/Riesz gases concerning above topics.
@@
We also emphasize that in the work of Lacker [34] and Lacker-Le Flem [35], it is shown that using the BBGKY hierarchy that the optimal local convergence rate of the relative entropy between $k$-marginals is $O(k/N)$ for systems with bounded interactions, while previous global arguments can only obtain the suboptimal $O(\sqrt{k/N})$ rate. This is verified by calculating the local relative entropy and solving the ODE hierarchy by iteration. Most recently in S. Wang [55] the optimal rate has been extended to the 2D Biot-Savart case with large viscosity on the torus also using the BBGKY hierarchy. Qualitative propagation of chaos for the 2D Vlasov-Poisson-Fokker-Planck equation has also been established by Bresch-Jabin-Soler [4] by hierarchy estimates. Recently, a new duality based method for the mean-field limit problem was introduced by Bresch-Duerinckx-Jabin [3], which requires only that the interacting kernel is in $L^{2}$.
@@ 1.4. Outline of the Article
The rest of this paper is organized as follows. We first present the proof of our main results in Section 2, acknowledging some logarithmic growth estimates of the limit density. The proof is divided into several subsections, respectively giving the time evolution of the relative entropy, some useful functional inequality lemmas and theorems, and the main controlling arguments. We shall provide the global and local relative entropy control respectively, and also set up an independent subsection for the proof of an ODE hierarchy proposition. In Section 3 we apply the results of Carlen-Loss and Osada and the parabolic maximum principle to give the needed regularity estimates for the limit density. In Section 4 we first present the Grigor’yan parabolic maximum principle as the key tool, and then provide our desired logarithmic growth estimates, i.e. growth controls on $\nabla\log g$ and $\nabla^{2}\log g$ under certain initial conditions.
@@ 2.1. Time Evolution of Relative Entropy
We present in this subsection the time evolution of the normalized relative entropy, which is computed directly using the evolution PDE of the two systems. First, we give the computations for the global relative entropy.
@@
$\displaystyle\frac{\,\mathrm{d}}{\,\mathrm{d}t}H_{N}(G^{N}|g^{N})=$ $\displaystyle-\frac{\sigma}{N}\sum_{i=1}^{N}\int_{\mathbb{D}^{N}}G^{N}\Big{|}\nabla_{x_{i}}\log\frac{G^{N}}{g^{N}}\Big{|}^{2}\,\mathrm{d}Z^{N}$
$\displaystyle-\frac{1}{N^{2}}\sum_{i,j=1}^{N}\int_{\mathbb{D}^{N}}G^{N}\Big{(}m_{j}K(x_{i}-x_{j})-K\ast\omega(x_{i})\Big{)}\cdot\nabla\log g(m_{i},x_{i})\,\mathrm{d}Z^{N}.$
@@
while the factorzied law $g^{N}$ solves (1.8):
@@
$\displaystyle$=\,= $\displaystyle\frac{1}{N}\int_{\mathbb{D}^{N}}\partial_{t}G^{N}(\log G^{N}+1-\log g^{N})+\frac{G^{N}}{g^{N}}\partial_{t}g^{N}\,\mathrm{d}Z^{N}$
$\displaystyle$=\,= $\displaystyle\frac{1}{N}\int_{\mathbb{D}^{N}}\Big{(}\sigma\sum_{i=1}^{N}\Delta_{x_{i}}G^{N}-\frac{1}{N}\sum_{i,j=1}^{N}m_{j}K(x_{i}-x_{j})\cdot\nabla_{x_{i}}G^{N}\Big{)}(\log G^{N}+1-\log g^{N})\,\mathrm{d}Z^{N}$
$\displaystyle+\frac{1}{N}\int_{\mathbb{D}^{N}}\Big{(}\sigma\sum_{i=1}^{N}\Delta_{x_{i}}g^{N}-\sum_{i=1}^{N}K\ast\omega(x_{i})\cdot\nabla_{x_{i}}g^{N}\Big{)}\frac{G^{N}}{g^{N}}\,\mathrm{d}Z^{N}.$
@@
$\displaystyle-\frac{\sigma}{N}\sum_{i=1}^{N}\int_{\mathbb{D}^{N}}\nabla_{x_{i}}G^{N}\cdot\nabla_{x_{i}}(\log G^{N}-\log g^{N})\,\mathrm{d}Z^{N}$
$\displaystyle+\frac{1}{N^{2}}\sum_{i,j=1}^{N}\int_{\mathbb{D}^{N}}m_{j}K(x_{i}-x_{j})\cdot\nabla_{x_{i}}(\log G^{N}-\log g^{N})G^{N}\,\mathrm{d}Z^{N}$
$\displaystyle-\frac{\sigma}{N}\sum_{i=1}^{N}\int_{\mathbb{D}^{N}}\nabla_{x_{i}}g^{N}\cdot\nabla_{x_{i}}\frac{G^{N}}{g^{N}}\,\mathrm{d}Z^{N}$
$\displaystyle-\frac{1}{N}\sum_{i=1}^{N}\int_{\mathbb{D}^{N}}\Big{(}K\ast\omega(x_{i})\cdot\nabla_{x_{i}}g^{N}\Big{)}G^{N}\,\mathrm{d}Z^{N}.$
@@
Now we turn to the computations for the local relative entropy.